# MyFramelinecolor=black, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=codeBackground

### v3

@@ Abstract
Our main result is a PoUW for the task of Matrix Multiplication $\mathsf{MatMul}(A,B)$ of arbitrary matrices with $1+o(1)$ multiplicative overhead compared to naïve $\mathsf{MatMul}$ (even in the presence of Fast Matrix Multiplication-style algorithms, which are currently impractical). We conjecture that our protocol has optimal security in the sense that a malicious prover cannot obtain any significant advantage over an honest prover. This conjecture is based on reducing hardness of our protocol to the task of solving a batch of low-rank random linear equations which is of independent interest.
@@ 1 Introduction
The concept of proofs of work (PoW), i.e., easy-to-check proofs of computational effort, was proposed in 1992 in a seminal work of Dwork and Naor [DN92]. Their original motivation was to discourage spam email messages by requiring the sender to compute a moderately hard, but not intractable, function in order for the email to go through. This idea is far more general than combating spam emails; indeed, it can be used as a method for limiting access to any resource.
[…]
Perhaps the most well-known application of Back’s PoW idea is in the Bitcoin cryptocurrency network [Nak08, GKL24]. Instead of using PoW to deter malicious email senders, Bitcoin’s PoW is used to enable competitive mining and thereby to secure the network. Specifically, a Bitcoin miner runs a program that collects unconfirmed transactions from the network to form a block. A block is only accepted by the network if its hash meets a certain difficulty target, (essentially) described by the number of trailing 0s needed in the hash output.
@@
Thus, an outstanding question in this context is whether it is necessary for the mining process to be so wasteful and have no other use beyond securing and driving the blockchain. A similar question was raised already in the 1992 work of Dwork and Naor, in the context of using PoW as a spam filter [DN92]:
@@
  • Economic value: The results of computations performed by miners should be of interest, and moreover, someone must be willing to pay for the result of the computation (formally, a task is $\kappa$-useful, if someone is willing to pay $\kappa\$$ for its completion).
@@
The first formal treatment of PoUW we are aware of is due to Ball et al. [BRSV17] wherein the authors suggest a proof of work mechanism (in the plain model) based on the conjectured hardness of various fine-grained algorithmic problems. They termed the proof of work useful because the miner could choose the instance of the task it wishes to solve by itself. Unfortunately, the overhead of the honest prover was significant, requiring the prover to solve poly-log many different instances of the task just to convince the verifier that a single instance was solved. In other words, their scheme did not satisfy the efficiency property from above and in a follow-up version of their manuscript [BRSV18] they retracted all claims about “usefulness.”
The absence of a PoUW (prior to this work) is not due to lack of trying. Not only was this problem mentioned by Dwork and Naor [DN92], PoUW has been the holy-grail of distributed consensus since the early days of Bitcoin, because it avoids the tradeoff between resource-efficiency and security, present in proof-of-stake (PoS) or other energy-efficient alternatives. As Vitalik Buterin, co-founder of Ethereum, articulates in his 2019 blog post: [^3] https://vitalik.eth.limo/general/2019/11/22/progress.html, Bullet 7.
@@
Even a theoretical construction of PoUW was elusive. Specifically, we are not aware of any suggestion for a truly useful proof of work mechanism where the miner (and not the network) gets to choose their own instance. We refer to a manuscript of Dotan and Tochner [DT20] surveying prior approaches, identifying for each one which properties of a “non-wasteful” and “truly useful” proof of work they do not satisfy.
@@ 1.1 Our Contributions
The discovery of Fast Matrix Multiplication (FMM) algorithms by Strassen [Str69] and subsequent improvements (starting with [CW87] and most recently [VXXZ24]), allowing to multiply $n\times n$ matrices in sub-cubic time $n^{\omega}\ll n^{3}$, had a profound theoretical impact in computer science and algorithm design. Our PoW protocol can be instantiated and applied with any of the above FMMs as the baseline algorithm. Unfortunately, FMM algorithms are unlikely to be practical on any imaginable hardware, as they are are highly asymptotic (earning them the infamous name “galactic algorithms"), or require significant memory.
Our protocol’s description and security analysis hold even in the presence of FMM algorithms, assuming that they are used in the protocol itself. Nevertheless, while reading the paper it is most convenient to consider the most practically-efficient naïve algorithm for $n\times k$ and $k\times m$, which runs in $\Theta(nkm)$ time. We remark that the essentially the only non-black-box property of naïve MatMul algorithm the protocol uses is self-reducibility: computing $A\cdot B$ by decomposing it into smaller $(n/r)\times(n/r)$ MatMul of $r\times r$ block matrices (storing intermediate values in the block-wise inner products). This self-reducibility property of MatMul is all the more used in FMM divide-and-conquer algorithms, hence if FMM becomes practical, it should be possible to adapt our protocol to the $n^{\omega}$ baseline as opposed to $n^{3}$.
@@ 1.2 Related Work
There have been various proposals to completely get rid of the wastefulness of PoW mining by requiring the “waste” of a different resource. Most well known is the notion of “proof of stake” where participants in the network need to lock up a sum of money and thereby receive voting rights proportion to their quantity of holdings (e.g., [BLMR14, GHM+17, KKKZ19] to mentioned just a few). Other resources, such as space [DFKP15, PKF+18, CP19], have been studied and deployed in systems. While these avoid energy waste, they incur waste in other domains, and they all introduce new economic and cryptographic challenges and tradeoffs due to the nature of the other resource being utilized to secure the network.
@@
In two recent beautiful works by Vaikuntanathan and Zamir [VZ25] and Braverman and Newman [BN25], the authors discovered a technique for sampling pseudorandom matrices such that whomever has a secret key can multiply them in near-linear time. The suggested constructions rely on standard cryptographic assumptions, i.e., learning parity with noise (LPN). The authors of [VZ25] suggest to use those so-called “trapdoor matrices” in various algorithmic applications that rely on products random matrices in order to obtain a significant speedup; among a series of applications, they emphasize a faster-than-known classification inference algorithm. The authors of [BN25] suggest applications of a similar technique in a context of secure delegation of linear algebra, wherein a (weak) client wishes to compute some public function of its private data. They achieve such protocols for linear public functions (i.e., matrix-matrix multipllication or matrix-vector multiplication) by “masking” the client’s private input with a “trapdoor matrix”.
@@ 2 Overview of our PoUW
Correctness of a PoUW says that $C=A\cdot B$ (with probability 1) and also $\mathsf{Verify}(\sigma,\pi)=1$ with probability $\epsilon$ (over the randomness of $\mathsf{Solve}$ and over a uniformly random $\sigma$), where $\epsilon>0$ is a parameter. Hardness of a PoUW says that no one can create an alternative $\mathsf{Solve}^{*}$ procedure which is able to run in time $t(n)\cdot(1+o(1))$ and succeed in creating $\pi$’s that are accepted with probability noticeably higher than $\epsilon$. That is, with the same computational budget there is no way to noticeably outperform the honest prover, $\mathsf{Solve}$.
We did not put any restriction on the running time of $\mathsf{Verify}$ in order to emphasize the challenging parts of the notion of a PoUW. Our final scheme will be quite efficient also in communication and verifier complexity; we discuss these later in the overview.
@@
Since $E$ and $F$ are uniformly random matrices, the matrix $C^{\prime}$ is uniformly random, no matter what $A$ and $B$ are. Therefore, $z$ will be below the above-mentioned threshold with probability exactly $\epsilon$, as needed. Suppose that $t(n)$ is the time spent by the prover on matrix multiplication in step 2. Note that $t(n)=O(n^{3})$ if we use classical matrix multiplication, and $t(n)=n^{\omega}$ if we use fast matrix multiplication, where the current record is $\omega<2.371339$ [ADV+24]. Besides step 2, all operations are linear time in the input size and are thus negligible. Overall, the prover’s runtime in the above system is $t(n)\cdot(1+o(1))$.
For the above system to be a PoW, we need to argue that there is no way to perform noticeably better than the honest prover. In our system, the verifier checks consistency of $z$ and so the only way to perform better is if an attacker could somehow come up with $A$ and $B$ for which it can compute $z$ faster than doing a generic matrix multiplication algorithm. Suppose otherwise, then we could use such a prover to compute the product of two random matrices in time $o(t(n))$ — a major breakthrough in algorithms. Assuming no such speedup is possible for random matrices, we conclude that a malicious prover must run in time at least $t(n)$, as much as the honest prover one up to additive lower-order terms.
@@
As we have already explained, the above scheme satisfies correctness in the sense that a verifier will accept an honestly generated proof with probability $\epsilon$ and also $C$ is equal to $A\cdot B$. For efficiency, notice that steps 1 and 4 can be implemented in time $O(n^{2}\cdot r)$ and step 3 requires hashing a string of length $(n/r)^{3}$ which takes time $o(n)$. Step 2 requires computing the whole transcript of the matrix multiplication (which includes the output, $C^{\prime}$). One way to compute it is directly running the classical matrix multiplication algorithm (in time $O(n^{3})$). However, there is a more efficient way using fast matrix multiplication techniques. Indeed, every intermediate $r\times r$ block can be computed via a fast rectangular matrix multiplication that can be done non-trivially in some regimes of parameters. For instance, one can multiply a $n\times n^{0.3}$ matrix with a $n^{0.3}\times n$ matrix in time $n^{2+o(1)}$ [VXXZ24]. Thus, in some regimes one can implement step 2 (and so step 3) in time $O(n^{3}/r)$. Depending on $r$, let us denote the current state of the art runtime of the above algorithm for step 2 by $t_{r}(n)$. Since the latter is the dominant term, we get that the total runtime of the honest prover is $t_{r}(n)\cdot(1+o(1))$.
@@ 3 Open Problems and Future Directions
Some natural and interesting candidates are Linear Equations, Graph search and routing problems, and database problems like Pattern-Matching and string problems (e.g., DNA sequencing). Some of these functions lack a natural decomposable structure into sub-problems $f(x)=\bigotimes g(x_{i})$ like MatMul, so it is less clear how to use our notion of random self-reducibility on subproblems.
@@ 4 Preliminaries
For an integer $n\in\mathbb{N}$, we denote $[n]=\{1,\ldots,n\}$. For a set $D$, $x\leftarrow D$ denotes sampling an element from $D$ uniformly at random. For a distribution $\mathcal{D}$ over a set $D$, $x\leftarrow\mathcal{D}$ denotes sampling an element of $D$ according to $\mathcal{D}$. By $U_{m}$ we denote the uniform distribution over $\{0,1\}^{m}$.
@@ 5 Definitional Framework for Proofs of Useful Work
Classical proofs of work [DN92] only require the last property to hold, and have no associated useful computation. Usually, one can leverage classical proofs of work to generate random instances of “useful” problems by sampling them from the proof of work challenge. In this way, for appropriate tasks (where average case complexity is well understood), a prover is guaranteed to spend a certain amount of work in producing proofs. However, while the algorithm being executed for generating a proof of work is related to a concrete computational task, the instance itself is random and detached from any fixed instance that someone may actually want to solve. Thus, such systems are not considered useful according to the above.
@@
It is very common to see definitions of proofs of work as non-interactive algorithms rather than interactive protocols. We may jump back and forth between the two methods, depending on the context. We remark that all of our protocols are public-coin and therefore we can use the Fiat-Shamir heuristic [FS86] to remove the interaction and prove security in the Random Oracle (RO) model . Thus, in what follows we assume the RO model, meaning that all parties have query access to a public random function.
@@
A PoUW for $f$ consists of two algorithms $\mathsf{PoUW}=(\mathsf{Solve},\mathsf{Verify})$. These algorithms must satisfy the properties $(\alpha,\beta)$-efficiency, $\epsilon$-completeness, hardness for functions $\alpha,\beta,\epsilon$, as described next:
$(\alpha,\beta)$-efficiency: For any $\sigma\in\{0,1\}^{\lambda}$ and any $\mathsf{x}\in\{0,1\}^{n}$:
  • $\mathsf{Solve}(\sigma,\mathsf{x})$ runs in time $\alpha(t(n))$.
  • $\mathsf{Verify}(\sigma,\pi)$ runs in time $\beta(t(n))$ for any possible output $\pi$ of $\mathsf{Solve}$.
$\epsilon$-completeness: For every $\lambda,n\in\mathbb{N}$, every $\sigma\in\{0,1\}^{\lambda}$, and every $\mathsf{x}\in\{0,1\}^{n}$, it holds that
@@
Hardness: For any $\delta>0$ and any algorithm $\mathsf{Solve}^{\star}$ with polynomial-time preprocessing and whose online running time is $(\alpha(t(n)))^{1-\delta}$, such that
$\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]<\epsilon(n),$
@@
In the above we give the adversary a smaller compute budget and assume that it cannot perform as well as the honest solver. There are many alternative ways to define hardness, depending on how flexible or strict one wants to be. For example, one could be more nuanced about the ratio between the work and the probability of verification succeeding. This is, in fact, more aligned with the typical way proofs of work are used in practice: if one invests twice more work, then its reward is factor two higher. The hardness property in this case can be formalized by parametrized the running time of the attacker by $\gamma(\alpha(n))$ and saying that $\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]\leq(\gamma(\alpha(n))/\alpha(n))\cdot\epsilon(n)$. For example, if the adversary invests half of the resources the honest solver does, then its expected reward is exactly half.
Additionally, one can introduce stronger notions of security, considering the success probability of a malicious prover when working on multiple instances in parallel (i.e., whether amortization exists), or when the adversary receives challenges in a streaming fashion and it only needs to show slight advantage over the naive process. The latter property is essentially what is needed for Bitcoin-style blockchains, and we formalize the security definition in Appendix B. We conjecture that our PoW mechanism will satisfy all properties satisfied by the PoW of Bitcoin, and focus in the main body on the simpler definition to simplify presentation and convey the main new ideas.
A PoUW scheme is said to be optimal if $\mathsf{Solve}$ introduces minimal overhead over directly solving $\mathsf{x}$, and if no adversary can obtain noticeable advantage over the honest prover. We also require non-trivial efficiency on the verifier. We make this explicit in the next definition.
Definition 5.2 (Optimal PoUW).
A PoUW for $f$, denoted $(\mathsf{Solve},\mathsf{Verify})$ is said to be optimal if
  • $\alpha(t(n))=t(n)\cdot(1+o(1))$ and $\beta(t(n))=t(n)\cdot(1-1/c)$ for $c>0$.
  • $\epsilon$ is any constant between 0 and 1.
We conclude this section with several remarks about the definition.
@@ A PoUW for $f$, denoted $(\mathsf{Solve},\mathsf{Verify})$ is said to be optimal if
We require that the (honest) prover outputs the correct $\mathsf{y}$ for each $\mathsf{x}$ it wants to solve (with probability 1). We do so for simplicity and because our construction will satisfy this strong requirement. However, it is plausible to consider a weaker definition where a correct solution would be outputted only with some reasonably high probability.
@@
We parameterize the efficiency of $\mathsf{Solve}$ using a general function $\alpha$. However, for practical purposes, we believe that the overhead of $\mathsf{Solve}$ should be as small as possible. Indeed, with constant multiplicative overhead, there is already a PoUW where solving the task and proving to the verifier that work has been done are done separately and independently (see Section 5.1 for details). Thus, the interesting regime is where $\alpha(t(n))\leq c\cdot t(n)$ where $c\ll 2$. In our protocol, we will achieve the best one can hope for $\alpha(t(n))\leq t(n)\cdot(1+o(1))$.
@@
One could impose additional constraints on the efficiency of the protocol, e.g., that the proof $\pi$ sent from $P$ to $V$ is very short, or that the verifier’s running time is sufficiently small. We will discuss these later in the paper.

### v4

# Arbitrary Matrix Multiplication

@@ Abstract
Our main result is a PoUW for the task of Matrix Multiplication $\mathsf{MatMul}(A,B)$ of arbitrary matrices with $1+o(1)$ multiplicative overhead compared to naïve $\mathsf{MatMul}$. We conjecture that our protocol has optimal security in the sense that a malicious prover cannot obtain any significant advantage over an honest prover. This conjecture is based on reducing hardness of our protocol to the task of solving a batch of low-rank random linear equations which is of independent interest.
@@ 1 Introduction
The concept of proofs of work (PoW), i.e., easy-to-check proofs of computational effort, was proposed in 1992 in a seminal work of Dwork and Naor [DBLP:conf/crypto/DworkN92]. Their original motivation was to discourage spam email messages by requiring the sender to compute a moderately hard, but not intractable, function in order for the email to go through. This idea is far more general than combating spam emails; indeed, it can be used as a method for limiting access to any resource.
A concretely efficient and very simple instantiation of a PoW (in the random oracle model) was proposed by Back in 1997 under the name Hashcash and formalized in a paper five years later [back2002hashcash]. Since then, the concept has been used as a denial-of-service counter measure technique in a number of systems. One way to describe the Hashcash algorithm’s idea is via a protocol between a prover and a verifier, wherein the verifier gets eventually convinced that the prover did a moderate amount of work. The protocol works with a cryptographic hash function $H$, such as SHA256 or SHA3. At a very high level, the verifier sends a random string $\sigma$ as a challenge and the goal of the prover is to find a string $x$ such that $H(\sigma\|x)$ starts with $t$ leading 0s. [^1] The symbol $\|$ represents string concatenation. Verification is very easy — just count leading 0s — and can even be done by a human eye. The soundness of the protocol relies on the fact that it is impossible to find an $x$ as above faster than doing a brute-force search, assuming $H$ is a random oracle.
Perhaps the most well-known application of Back’s PoW idea is in the Bitcoin cryptocurrency network [nakamoto2008bitcoin, GarayKL24]. Instead of using PoW to deter malicious email senders, Bitcoin’s PoW is used to enable competitive mining and thereby to secure the network. Specifically, a Bitcoin miner runs a program that collects unconfirmed transactions from the network to form a block. A block is only accepted by the network if its hash meets a certain difficulty target, (essentially) described by the number of trailing 0s needed in the hash output.
@@
Thus, an outstanding question in this context is whether it is necessary for the mining process to be so wasteful and have no other use beyond securing and driving the blockchain. A similar question was raised already in the 1992 work of Dwork and Naor, in the context of using PoW as a spam filter [DBLP:conf/crypto/DworkN92]:
@@
  • Economic value: The results of computations performed by miners should be of interest, and moreover, someone must be willing to pay for the result of the computation (formally, a task is $\kappa$-useful, if someone is willing to pay $\kappa\mathdollar$ for its completion).
@@
The first formal treatment of PoUW we are aware of is due to Ball et al. [BallRSV17a] wherein the authors suggest a proof of work mechanism (in the plain model) based on the conjectured hardness of various fine-grained algorithmic problems. They termed the proof of work useful because the miner could choose the instance of the task it wishes to solve by itself. Unfortunately, the overhead of the honest prover was significant, requiring the prover to solve poly-log many different instances of the task just to convince the verifier that a single instance was solved. In other words, their scheme did not satisfy the efficiency property from above and in a follow-up version of their manuscript [DBLP:conf/crypto/BallRSV18] they retracted all claims about “usefulness.”
The absence of a PoUW (prior to this work) is not due to lack of trying. Not only was this problem mentioned by Dwork and Naor [DBLP:conf/crypto/DworkN92], PoUW has been the holy-grail of distributed consensus since the early days of Bitcoin. Indeed, it avoids the tradeoff between resource-efficiency and security, present in proof-of-stake (PoS) or other energy-efficient alternatives. As Vitalik Buterin, co-founder of Ethereum, articulates in his 2019 blog post: [^3] https://vitalik.eth.limo/general/2019/11/22/progress.html, Bullet 7.
@@
Even a theoretical construction of PoUW was elusive. Specifically, we are not aware of any suggestion for a truly useful proof of work mechanism where the miner (and not the network) gets to choose their own instance. We refer to a manuscript of Dotan and Tochner [DBLP:journals/corr/DotanTochner] surveying prior approaches, identifying for each one which properties of a “non-wasteful” and “truly useful” proof of work they do not satisfy.
@@ 1.1 Our Contributions
The discovery of Fast Matrix Multiplication (FMM) algorithms by Strassen [strassen1969gaussian] and subsequent improvements (starting with [coppersmith1987matrix] and most recently [DBLP:conf/soda/WilliamsXXZ24]), allowing to multiply $n\times n$ matrices in sub-cubic time $n^{\omega}\ll n^{3}$, had a profound theoretical impact in computer science and algorithm design. Our PoW protocol can be instantiated and applied with any of the above FMMs as the baseline algorithm. Unfortunately, FMM algorithms are unlikely to be practical on any imaginable hardware, as they are are highly asymptotic (earning them the infamous name “galactic algorithms"), or require significant memory.
Indeed, since our work is meant to serve as a practical alternative to Bitcoin, our analysis assumes naïve MatMul as the most practically-efficient baseline algorithm for $n\times k$ and $k\times m$, which runs in $\Theta(nkm)$ time.
We stress that our protocol’s security analysis holds even in the presence of FMM algorithms, assuming that they are used to attack the protocol itself. Developing an analogues PoUW scheme to cuPOW  for Strassen-like FMM algorithms is an intriguing question from a theoretical standpoint, but this is not the focus of this paper.
@@ 1.2 Related Work
There have been various proposals to completely get rid of the wastefulness of PoW mining by requiring the “waste” of a different resource. Most well known is the notion of “proof of stake” where participants in the network need to lock up a sum of money and thereby receive voting rights proportion to their quantity of holdings (e.g., [DBLP:journals/sigmetrics/BentovLMR14, DBLP:journals/iacr/GiladHMVZ17, DBLP:conf/sp/KerberKKZ19] to mentioned just a few). Other resources, such as space [DBLP:conf/crypto/DziembowskiFKP15, DBLP:conf/fc/ParkKFGAP18, cohen2019chia], have been studied and deployed in systems. While these avoid energy waste, they incur waste in other domains, and they all introduce new economic and cryptographic challenges and tradeoffs due to the nature of the other resource being utilized to secure the network.
@@
In two recent beautiful works by Vaikuntanathan and Zamir [VaikuntanathanZamir25] and Braverman and Newman [BravermanNewman25], the authors discovered a technique for sampling pseudorandom matrices such that whomever has a secret key can multiply them in near-linear time. The suggested constructions rely on standard cryptographic assumptions, i.e., learning parity with noise (LPN). The authors of [VaikuntanathanZamir25] suggest to use those so-called “trapdoor matrices” in various algorithmic applications that rely on products of random matrices in order to obtain a significant speedup; among a series of applications, they emphasize a faster-than-known classification inference algorithm. The authors of [BravermanNewman25] suggest applications of a similar technique in a context of secure delegation of linear algebra, wherein a (weak) client wishes to compute some public function of its private data. They achieve such protocols for linear public functions (i.e., matrix-matrix multipllication or matrix-vector multiplication) by “masking” the client’s private input with a “trapdoor matrix”.
@@ 2 Overview of our PoUW
Correctness of a PoUW says that $C=A\cdot B$ (with probability 1) and also $\mathsf{Verify}(\sigma,\pi)=1$ with probability $\epsilon$ (over the randomness of $\mathsf{Solve}$ and over a uniformly random $\sigma$), where $\epsilon>0$ is a parameter. Hardness of a PoUW says that no one can create a procedure that runs in time $t(n)\cdot(1+o(1))$ and succeed in creating $\pi$’s that are accepted with probability noticeably higher than $\epsilon$ (notice that this procedure does not need to even multiply matrices). That is, with the same computational budget there is no way to noticeably outperform the honest prover, $\mathsf{Solve}$.
We did not put any non-trivial restriction on the running time of $\mathsf{Verify}$ in order to emphasize the challenging parts of the notion of a PoUW. Our final scheme will be quite efficient also in communication and verifier complexity; we discuss this later.
@@
Since $E$ and $F$ are uniformly random matrices, the matrix $C^{\prime}$ is uniformly random, no matter what $A$ and $B$ are. Therefore, $z$ will be below the above-mentioned threshold with probability exactly $\epsilon$, as needed. Suppose that $t(n)$ is the time spent by the prover on matrix multiplication in step 2. Note that $t(n)=O(n^{3})$ if we use classical matrix multiplication, and $t(n)=n^{\omega}$ if we use fast matrix multiplication, where the current record is $\omega<2.371339$ [abs-2404-16349]. Besides step 2, all operations are linear time in the input size and are thus negligible. Overall, the prover’s runtime in the above system is $t(n)\cdot(1+o(1))$.
For the above system to be a PoW, we need to argue that there is no way to perform noticeably better than the honest prover. In our system, the verifier checks consistency of $z$ and so the only way to perform better is if an attacker could somehow come up with $A$ and $B$ for which it can compute $z$ faster than doing a generic matrix multiplication algorithm. Suppose otherwise: we could use such a prover to compute the product of two random matrices in time $o(t(n))$ — a major breakthrough in algorithms. Assuming no such speedup is possible for random matrices, we conclude that a malicious prover must run in time at least $t(n)$, as much as the honest prover one up to additive lower-order terms.
@@
As we have already explained, the above scheme satisfies correctness in the sense that a verifier will accept an honestly generated proof with probability $\epsilon$ and also $C$ is equal to $A\cdot B$. For efficiency, notice that steps 1 and 4 can be implemented in time $O(n^{2}\cdot r)$ and step 3 requires hashing a string of length $(n/r)^{3}$ which takes time $o(n)$. Step 2 requires computing the whole transcript of the matrix multiplication (which includes the output, $C^{\prime}$). One way to compute it is directly running the classical matrix multiplication algorithm (in time $O(n^{3})$). However, there is a more efficient way using fast matrix multiplication techniques. Indeed, every intermediate $r\times r$ block can be computed via a fast rectangular matrix multiplication that can be done non-trivially in some regimes of parameters. For instance, one can multiply a $n\times n^{0.3}$ matrix with a $n^{0.3}\times n$ matrix in time $n^{2+o(1)}$ [DBLP:conf/soda/WilliamsXXZ24]. Thus, in some regimes one can implement step 2 (and so step 3) in time $O(n^{3}/r)$. Depending on $r$, let us denote the current state of the art runtime of the above algorithm for step 2 by $t_{r}(n)$. Since the latter is the dominant term, we get that the total runtime of the honest prover is $t_{r}(n)\cdot(1+o(1))$.
@@
Remark 2.2 (Improving verifier’s complexity).
The scheme as described above is relatively expensive on the verifier’s side. Indeed, without any optimizations, the verifier needs to repeat the prover’s work to get convinced that $\pi$ is “valid”. Using standard tools we can improve this. Indeed, we can delegate the verifier’s work (when $\mathsf{Verify}$ should accept) to the prover by asking the latter to compute a SNARK certifying that the verifier would have accepted the proof had it seen it. Furthermore, if privacy of $A,B$ is an issue, the prover can use a zkSNARK.
Remark 2.3 (Using SNARKs as a source of hardness?).
The above paragraph leads to a natural idea of using the SNARK computation itself as a source of hardness. Indeed, we could “force” the prover to follow a specific algorithm by asking it to generate a SNARK proof attesting to this fact. For this purpose one could avoid noising and denoising altogether if the SNARK computation itself is keyed (as in Micali’s [DBLP:conf/focs/Micali94] random oracle-based SNARK). This indeed forces the prover to work and prevents a malicious prover from doing shortcuts. The problem is that we typically think of $\epsilon$, the probability that $\mathsf{Solve}$ outputs an accepting proof, as being small. Nevertheless, the above approach requires the prover to compute a SNARK every time, making the efficiency property of the scheme quite poor. (SNARKs require significant time to compute.) In our suggestion, on the other hand, only when the (rare) event that a valid $\pi$ is found, then we expect the solver to generate the (zk)SNARK, and so the cost of this computation is amortized and becomes negligible.
@@ 3 Open Problems and Future Directions
Some natural and interesting candidates are solving linear equations, graph search and routing problems, and database problems like pattern matching and string problems (e.g., DNA sequencing). Some of these functions lack a natural decomposable structure into sub-problems $f(x)=\bigotimes g(x_{i})$ like MatMul, so it is less clear how to use our notion of random self-reducibility on subproblems.
@@
A provably secure scheme.
A fascinating question is whether one can design a PoUW scheme for matrix multiplication (or other “useful” operations) with provable security under more standard (albeit possibly strong) assumptions. For instance, is there a PoUW from general purpose program obfuscation assumptions?
Problem 3.3.
Is there a PoUW scheme for useful operations from more standard or well-studied assumptions?
@@ 4 Preliminaries
For an integer $n\in\mathbb{N}$, we denote $n]=\{1,\ldots,n\}$. For a set $D$, $x\leftarrow D$ denotes sampling an element from $D$ uniformly at random. For a distribution $\mathcal{D}$ over a set $D$, $x\leftarrow\mathcal{D}$ denotes sampling an element of $D$ according to $\mathcal{D}$. By $U_{m}$ we denote the uniform distribution over $\{0,1\}^{m}$.
@@ 5 Definitional Framework for Proofs of Useful Work
Classical proofs of work [DBLP:conf/crypto/DworkN92] only require the last property to hold, and have no associated useful computation. Usually, one can leverage classical proofs of work to generate random instances of “useful” problems by sampling them from the proof of work challenge. In this way, for appropriate tasks (where average case complexity is well understood), a prover is guaranteed to spend a certain amount of work in producing proofs. However, while the algorithm being executed for generating a proof of work is related to a concrete computational task, the instance itself is random and detached from any fixed instance that someone may actually want to solve. Thus, such systems are not considered useful according to the above.
@@
It is very common to see definitions of proofs of work as non-interactive algorithms rather than interactive protocols. We may jump back and forth between the two methods, depending on the context. We remark that all of our protocols are public-coin and therefore we can use the Fiat-Shamir heuristic [FiatS86] to remove the interaction and prove security in the Random Oracle (RO) model . Thus, in what follows we assume the RO model, meaning that all parties have query access to a public random function.
@@
An $\epsilon$-difficulty and $\alpha$-overhead PoUW for $f$ consists of two algorithms $\mathsf{Solve}$ and $\mathsf{Verify}$, satisfying the efficiency, completeness, and hardness properties, described next.
Efficiency: For any $\sigma\in\{0,1\}^{\lambda}$ and any $\mathsf{x}\in\{0,1\}^{n}$:
  • $\mathsf{Solve}(\sigma,\mathsf{x})$ runs in time $t(n)\cdot(1+\alpha(n))$.
  • $\mathsf{Verify}(\sigma,\pi)$ runs in time $c\cdot t(n)$ for constant $c\in(0,1)$ for any possible output $\pi$ of $\mathsf{Solve}$.
Completeness: For every $\lambda,n\in\mathbb{N}$, every $\sigma\in\{0,1\}^{\lambda}$, and every $\mathsf{x}\in\{0,1\}^{n}$, it holds that
@@
Hardness: There exists a constant $C>0$ such that for any algorithm $\mathsf{Solve}^{\star}$ running in time $t^{\prime}=t(n)$ with $t^{\prime}(n)\leq t(n)$ with polynomial time preprocessing, it holds that
$\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]<\max\left\{C\cdot\frac{t^{\prime}\cdot\epsilon(n)}{t},2^{-\lambda}\right\},$
@@
One can introduce stronger notions of security, considering the success probability of a malicious prover when working on multiple instances in parallel (i.e., whether amortization exists), or when the adversary receives challenges in a streaming fashion and it only needs to show slight advantage over the naive process. The latter property is essentially what is needed for Bitcoin-style blockchains, and we formalize the security definition in Appendix B. We conjecture that our PoW mechanism will satisfy all properties satisfied by the PoW of Bitcoin, and focus in the main body on the simpler definition to simplify presentation and convey the main new ideas. We conclude this section with several remarks about the definition.
@@
We require that the (honest) prover outputs the correct $\mathsf{y}$ for each $\mathsf{x}$ it wants to solve (with probability 1). We do so for simplicity and because our construction will satisfy this strong requirement. However, it is plausible to consider weaker definition where, either a correct solution or an approximation thereof is outputted with reasonably high probability.
@@
The efficiency of $\mathsf{Solve}$ is parametrized via a function $\alpha$ to which we refer as overhead. Obviously, the overhead should be as small as possible, so as to incentivize running $\mathsf{Solve}$ instead of solving $f$ and trying to find $\pi$ separately. Indeed, the latter already gives a PoUW with constant multiplicative overhead (that depends on $\epsilon$). See Section 5.1 for details. Thus, the interesting regime is where $\alpha(t(n))\ll t(n)$ and ideally $\alpha(n)=o(t(n))$. The latter is achieved by our protocol.
@@
One could impose additional constraints on the efficiency of the protocol, e.g., that the proof $\pi$ sent from $P$ to $V$ is very short, or that the verifier’s running time is sufficiently small. As we elaborate later in the paper, one can generically achieve these properties by using off-the-shelve tools such as succinct non-interactive arguments (SNARGs). (Note that computing the SNARG is only done when a “valid” $\pi$ is found, and so the complexity of computing the SNARG is amortized away.)